OpenStudy (anonymous):

How do I convert 115°26'27'' into a decimal?

jimthompson5910 (jim_thompson5910):

x degrees, y minutes, z seconds = x + y/60 + z/3600 115°26'27'' = 115 + 26/60 + 27/3600 115°26'27'' = ??

OpenStudy (anonymous):

I must be doing something wrong. I used your method and I got 115.4408, but my book says the answer is 262.2586.

jimthompson5910 (jim_thompson5910):

how is that even possible...the answer should be near 115

OpenStudy (anonymous):

That's what I thought! It's been driving me insane all day.

jimthompson5910 (jim_thompson5910):

the 26 minutes and 27 seconds are basically fractions of a whole degree, there's no way they turn 115 into 262

jimthompson5910 (jim_thompson5910):

either there's a typo, or there's been a problem mixup

jimthompson5910 (jim_thompson5910):

I'd just let it go for now and ask your teacher later

OpenStudy (anonymous):

Okay, thank you :)

jimthompson5910 (jim_thompson5910):

yw